# SummExecEdit: A Factual Consistency Benchmark in Summarization with Executable Edits

### v1

@@ Abstract
Detecting factual inconsistencies in summarization is critical, yet existing benchmarks lack the necessary challenge and interpretability for robust evaluation. In this paper, we introduce SummExecEdit, a novel benchmark leveraging executable edits to assess models on their ability to both detect factual errors and provide accurate explanations. The top-performing model, Claude3-Opus, achieves a joint detection and explanation score of only 0.49 in our benchmark, with individual scores of 0.67 for detection and 0.73 for explanation. Furthermore, we identify four primary types of explanation errors, with 45.4% of errors focusing on completely unrelated parts of the summary.
@@ 1 Introduction
Large language models (LLMs) have shown remarkable performance across various tasks by generating coherent responses while facing a major challenge with factual hallucination Huang et al. (2024). Various evaluation methods Laban et al. (2022); Fabbri et al. (2021, 2022); Tang et al. (2023); Luo et al. (2023); Yang et al. (2024) and inconsistency detection benchmarks have been proposed, and they used approaches such as editing ground truth texts to create intentional inconsistencies Laban et al. (2023) or generating benchmarks with LLMs where they label the factual accuracy of target texts post-generation. Tang et al. (2024b)
@@
However, previous approaches usually have broad and sweeping edits, i.e., trivial edits and multiple edits in one sample. Kim et al. (2024) in their work chose to entirely rely on humans to generate edits, because initial experiments with LLM-based edits yielded mostly trivial edits. Such edits make factual inconsistency errors easier to be detected by LLMs, for example, GPT4 achieves 82.4% detection accuracy on a previous inconsistency benchmark - SummEdits Laban et al. (2023). This shows that we need mechanisms to generate more complex edits. Moreover, high quality explanations about inconsistency in these detection benchmarks are missing, and they are necessary to evaluate the model’s ability to explain their detection result and reason over facts.
@@
To address these drawbacks, we propose to leverage the concept of executable edits Laban et al. (2024) to generate a challenging and interpretable benchmark called SummExecEdit - extending SummEdits Laban et al. (2023)  [^1] Our data is released at https://huggingface.co/datasets/Salesforce/summexecedit. While normal editing rewrites the entire summary with edits already applied, executable editing focuses on isolating and replacing a specific substring in the text showing what has changed, allowing precise changes that can introduce factual inconsistencies. Such edits helps LLM to better concentrate and come up with a complex yet granular and controlled edit with a meaningful explanation for same in a structured way, helping to eliminate broad and sweeping multiple edits, as shown in Figure 1.
@@
We compare executable edits to existing editing and show that executable edits are superior, providing in range of 18-25% of higher number of challenging samples with better explanations. We evaluate a wide range of open-source and open-api access LLMs on SummExecEdit and find that most of these models lack on the combined task of detection and explanation of error, with the best model - Claude3-Opus achieving a joint score of only 0.49, with individual scores of 0.67 for detection and 0.73 for explanation. We also perform error analyses for incorrect explanations generated by candidate LLMs and categorize them into four reasons, with 45.4% of errors focusing on completely unrelated parts of the summary.
2 Related Works
Editing and Annotation in Previous Benchmarks. Several previous works have assessed various metrics in relation to human judgments associated with hallucination and factual inconsistencies. Works prior to advances in language models focused on human and expert level Fabbri et al. (2021); Falke et al. (2019) while the very recent works focused on generating annotations with the help of language models Chhikara et al. (2024); Tang et al. (2024b); Laban et al. (2023). Although these models generate coherent texts with a good level of inter-annotator agreements with respect to consistency Laban et al. (2023), such benchmarks may struggle to keep the changes in the original data to minimal Nguyen et al. (2024). Executable edits with help of language models was first explored by Laban et al. (2024), where it focused on replacing a substring for document editing. We build upon the works of SummEdits Laban et al. (2023) to use executable edits to generate new factual consistency benchmark.
Detecting Factual Inconsistencies. Since automatic evaluation metrics such as ROUGE Lin (2004) or BLEU Papineni et al. (2002) cannot reliably detect factual inconsistencies, several non-LLM based methods have been developed such as question-based Fabbri et al. (2022) and entailment-based Laban et al. (2022); Zha et al. (2023); Tang et al. (2024a). In various recent works, LLMs have shown superior performance when used as a judge to evaluate factual consistencies Luo et al. (2023). We explore this further in various settings and also re-evaluate the modern-day LLMs on our benchmark generated using executable edits.
3 Executable Edits
Figure 1: Example comparison between non-executable and executable edit.
Moreover, this minimizes synthetic data in the benchmarks since a majority of the original data remains same. Executable edits can also have a wide usage for data augmentation.
@@ 4 SummExecEdit Benchmark
We leverage Laban et al. (2023) to build our benchmark across 10 domains such as News, Podcast, Bill, Sales calls, etc. Based on Table 1, we use both GPT4-Turbo and Claude3-Opus models to generate our benchmark using the executable prompt. Both the models are asked to generate six edits for each (document, summary) pair.
After generating these edits, we remove the trivial edits with help of GPT4-Turbo. We ask GPT4-Turbo to classify an edit as a date change, number change, antonym change, or others. Any edit classified as date, number, or antonym change is removed from the benchmark.
The final benchmark results in 2,121 factually inconsistent summaries. To balance out the entire benchmark, we use all 2,120 factually consistent edits from SummEdits benchmark resulting in a total of 4,241 samples in the final benchmark. Each of the 10 domains provide around 200-300 inconsistent summaries. The distribution of each domains is given in Table 6.
@@ 5 Results
To evaluate LLMs on SummExecEdit benchmark, we use two types of prompts. D&E - Detect and Explain error. Models need to detect if there is any factual inconsistency in summary, if yes, explain the inconsistency (Appendix F.3). E|D - Explain error given Detection. Given that the summary is inconsistent, models need to explain the inconsistency in the summary (Appendix F.4).
Table 2 provides the Detection Accuracy (DA) of all the models using prompt D&E. The best performing model on SummExecEdit , GPT4o, provides an accuracy of only around 73%. The overall results show that many LLMs struggle in detecting the factual error. As a reference, we also evaluate two non-LLM based approaches which use far lesser compute - AlignScore Zha et al. (2023) receiving 57.4% and MiniCheck Tang et al. (2024a) receiving 60.0% accuracy, which are better than several open-source LLMs.
5.1 Evaluating Explanations
The correlations for four prompts with different LLMs with respect to our manually annotated explanations are provided in appendix in Table 4. Selected explanations for manual annotations were shuffled and randomly selected. The model and either of the two prompts that generated those explanations were anonymized. The IAA between both the annotators are - Correlation of 0.885 and Cohen Kappa of 0.81. We think that the reference explanations generated at the time of edit itself are the best and of high quality, scoring 0.95 for 40 samples. The prompt EvalV4 works the best which suggests that evaluating explanations or reasoning of models, works better when we have access to what the edit/reason is or a reference explanation.
We use prompt EvalV4 with GPT4o model to give the Explanation Score (ES) of all models on 2,121 inconsistent summaries generated using both our prompts - D&E and E|D. The results are provided in Table 2.
@@ 5.1 Evaluating Explanations
Table 2: Scores for D&E (Prompt Detect and Explain error) and E|D (Prompt Explain error given Detection) for inconsistent summaries. DA - Detection Accuracy, DS - Detection Score, ES - Explanation Score, JS - Joint Score. ES evaluated by EvalV4 prompt using GPT4o. DA based on entire benchmark. DS, ES, and JS based only on inconsistent summaries. Last 5 models are open-source.
@@ 6 Conclusion
In this work, we first explored the executable editing with LLMs for generating benchmarks. Through our experiments, we show its superiority in generating a challenging benchmark and use it to generate a new factual consistency detection and explanation benchmark - SummExecEdit . We evaluate various LLMs over this benchmark for both detection and explanation of factual inconsistencies. The best performing model for detection - GPT4o achieves an accuracy of 73% while Claude3-Opus performs best on combined task of detection and explanation achieving a joint score of 0.49. These scores suggest that SummExecEdit is a challenging benchmark. We explore various prompt formats that lead to better auto-evaluation by LLMs and perform error-analysis for incorrect explanations generated by LLMs. We hope researchers and practitioners use executable edits to generate and augment data, and SummExecEdit can be used by LLM Developers for evaluating LLMs’ abilities to detect factual errors and reason about facts.
@@
While executable edits help in creating more challenging benchmark, it is important to note a few limitations of these edits and this work in general. First, generating benchmarks with executable edits require availability of ground truth (source, target) pairs which might not always be the case. Second, non-executable edits are equally capable of generating controlled edits but these are simple and low quality. If the end-goal is to only generate controlled edits and complexity does not matter, non-executable edits could be used as well. Last, we have only experimented using these edits for creating a single benchmark related to summarization. Further experiments in other domains, as well as for data augmentations, are required.
@@ Appendix A Model Details
Google Models. We used four Google models in our experiments - Gemini-1.5-Pro, Gemini-1.5-Flash, Gemini-Pro, and Palm-Bison. We used Google Cloud Platform to access Google models, which relies on the VertexAI API.
@@ Appendix B License Details and Terms of Use
Appendix C Annotation Tool
Appendix D Manual Annotation and Inter-Annotator Agreement
@@
Table 4 provides the Correlation and Balanced Accuracy for manual annotation of explanations based on EvalV1, EvalV2, EvalV3, and EvalV4 for four models under consideration.
Correlation BAcc
Model V1 V2 V3 V4 V1 V2 V3 V4
claude3-haiku 0.076 0.478 0.732 0.725 0.385 0.49 0.655 0.618
claude3-opus 0.484 0.782 0.801 0.774 0.46 0.76 0.776 0.738
gpt3.5-turbo 0.103 0.118 0.325 0.563 0.375 0.401 0.495 0.625
gpt4o 0.603 0.816 0.804 0.833 0.533 0.792 0.794 0.811
@@
Appendix E SummExecEdit Statistics
Table 6 provides statistics for different domains present in the DEFES benchmark.
Domain N %Inconsistent
SciTLDR 307 58.31
News 567 47.8
Podcast 344 58.72
BillSum 608 44.9
SamSum 450 51.11
Shakespeare 511 41.3
QMSum 334 50.3
ECTSum 416 47.12
Sales Email 368 57.1
Sales Call 336 53.87
Total 4241 50
Table 6: Statistics of the domains in SummExecEdit

### v2

@@ Abstract
Detecting factual inconsistencies in summarization is critical, yet existing benchmarks lack the necessary challenge and interpretability for robust evaluation. In this paper, we introduce SummExecEdit, a novel pipeline and benchmark leveraging executable edits to assess models on their ability to both detect factual errors and provide accurate explanations. The top-performing model, Claude3-Opus, achieves a joint detection and explanation score of only 0.49 in our benchmark, with individual scores of 0.67 for detection and 0.73 for explanation. We conduct detailed evaluations to assess the current state of models in this field and find that more than half of the 20+ LLMs in our study struggle with over 30% of the SummExecEdit benchmark. Additionally, we identify four primary types of explanation errors, with 45.4% of them involving a focus on completely unrelated parts of the summary.
@@ 1 Introduction
Large language models (LLMs) have shown remarkable performance across various tasks by generating coherent responses while facing a major challenge with factual hallucination Huang et al. (2024). Various evaluation methods Laban et al. (2022); Fabbri et al. (2021, 2022); Tang et al. (2023); Luo et al. (2023); Yang et al. (2024) and inconsistency detection benchmarks have been proposed, which used approaches such as editing ground truth texts to create intentional inconsistencies Laban et al. (2023), or generating benchmarks with LLMs where they label the factual accuracy of target texts post-generation Tang et al. (2024b).
@@
However, previous approaches often involve broad and sweeping edits, i.e., trivial and multiple edits within a single sample Laban et al. (2023). Kim et al. (2024) relied entirely on humans to generate edits in their work because initial experiments with LLM-based edits primarily produced trivial changes. LLM-generated multiple trivial edits make factual inconsistency errors easier for LLMs to detect. For instance, GPT-4 achieves an 82.4% detection accuracy on the SummEdits inconsistency benchmark Laban et al. (2023). However, relying on humans to generate edits and benchmarks is neither scalable nor cost-efficient. Therefore, we need mechanisms to generate more complex edits in a scalable and cost-effective manner. Moreover, high-quality explanations for inconsistencies in these detection benchmarks are missing. Such explanations are essential for evaluating a model’s ability to justify its detection results and reason over factual inconsistencies.
@@
To address these drawbacks, we propose to leverage the concept of executable edits Laban et al. (2024) to generate a challenging and interpretable benchmark called SummExecEdit - extending SummEdits  [^1] Our data is released at https://huggingface.co/datasets/Salesforce/summexecedit. While normal editing rewrites the entire summary with edits already applied, executable editing focuses on isolating and replacing a substring in the original summary, allowing precise changes that can introduce factual inconsistencies. Such edits helps LLM to better concentrate and come up with a complex yet granular and controlled edit with a meaningful explanation for the same in a structured way, helping to eliminate broad and sweeping multiple edits.
Figure 1: Overall contribution of our work. We start by first verifying and proving the superiority of executable editing over directly editing a summary. We use human annotation to label four different parameters. We then use SummExecEdit pipeline to create the full benchmark ensuring quality is not hampered when we scale up from our POC. We finally evaluate variety of LLMs on this benchmark using two different prompt types. For simplicity, we have not shown the minor details. Pipelines like SummExecEdit can be further used for creation of such datasets and benchmarks.
Example of SummExecEdit Original Summary The document depicts the banishment of Coriolanus, deemed an enemy to the people and his country. Despite the sadness of his family and friends, Coriolanus departs, promising to make them proud and maintain contact while they hope for a chance to repeal his banishment. Generating inconsistent summary with SummExecEdit Text to Replace "deemed an enemy to the people and his country" Replaced Text "deemed a traitor to the people and his country" Edited Summary generated with help of Executable Prompt The document depicts the banishment of Coriolanus, deemed an enemy to the people and his country deemed a traitor to the people and his country. Despite the sadness of his family and friends, Coriolanus departs, promising to make them proud and maintain contact while they hope for a chance to repeal his banishment. Explanation The document states that Coriolanus is deemed an enemy, but the summary escalates this to calling him a traitor, which has a more severe connotation not supported by the document. Inconsistent summary from SummEdits benchmark The document depicts the banishment of Coriolanus, deemed an enemy to his people and his country but a friend to his country. Despite the sadness of his family and friends foes, Coriolanus departs, promising to make them proud and maintain contact sever contact while they hope for a chance to repeal his banishment.
Figure 2: An example of comparison of inconsistent summaries. For SummExecEdit summary, LLM generates the text that should be replaced in the original summary with an alternative inconsistent text and an explanation for the inconsistency. The edit is controlled and granular - difficult to identify but still meaningful enough. An inconsistent summary directly generated with a direct prompt makes multiple trivial edits that are easy to find.
@@
Through our work, we aim to introduce an executable editing style for creating benchmarks and datasets, present a new benchmark - SummExecEdit for evaluating LLMs in hallucination detection and explanation within real-world systems, and thus identify the current challenges LLMs face in this field. We compare executable edits to normal direct editing and show that executable edits are superior, providing in range of 18-25% of higher number of challenging samples with better explanations. We evaluate a wide range of open-source and open-api access LLMs on SummExecEdit and find that most of these models lack on the combined task of detection and explanation of error, with the best model - Claude3-Opus achieving a joint score of only 0.49, with individual scores of 0.67 for detection and 0.73 for explanation. We find 60 inconsistent summaries being detected as consistent by all 20+ LLMs in our study and analyze their categories. We also perform error analyses for incorrect explanations generated by candidate LLMs and categorize them into four reasons, with 45.4% of errors focusing on completely unrelated parts of the summary. Overall contribution of our work is as shown in Figure 1.
@@
2 Related Works
Methods employed in creating previous benchmarks. Several previous works have assessed various metrics in relation to human judgments associated with hallucination and factual inconsistencies. Works prior to advances in language models focused on human and expert level Fabbri et al. (2021); Falke et al. (2019) while the very recent works focused on generating annotations with the help of language models Chhikara et al. (2024); Tang et al. (2024b); Laban et al. (2023). Although these models generate coherent texts with a good level of inter-annotator agreements with respect to consistency Laban et al. (2023), such benchmarks may struggle to keep the changes in the original data to minimal Nguyen et al. (2024). Executable edits with help of language models was first explored by Laban et al. (2024), where it focused on replacing a substring for document editing. We build upon the works of SummEdits Laban et al. (2023) to use executable edits to generate new factual consistency benchmark.
Detecting Factual Inconsistencies. Since automatic evaluation metrics such as ROUGE Lin (2004) or BLEU Papineni et al. (2002) cannot reliably detect factual inconsistencies, several non-LLM based methods have been developed such as question-based Fabbri et al. (2022) and entailment-based Laban et al. (2022); Zha et al. (2023); Tang et al. (2024a). Gekhman et al. (2023) introduced TrueTeacher - a method for generating synthetic data by annotating diverse model-generated summaries using a LLM that can enhance the performance of a state-of-the-art model. In various recent works, LLMs have shown superior performance when used as a judge to evaluate factual consistencies Luo et al. (2023). We explore this further in various settings and also re-evaluate modern-day LLMs on our benchmark generated using executable edits.
3 Executable Edits
Moreover, this minimizes synthetic data in the benchmarks since a majority of the original data remains same. An example of SummExecEdit is given in Figure 2
@@ 4 SummExecEdit Benchmark
We leverage Laban et al. (2023) to build our benchmark across 10 domains such as News, Podcast, Bills, Sales calls, etc. Based on Table 1, we use both GPT4-Turbo and Claude3-Opus models to generate our benchmark using the executable prompt. Both the models are asked to generate six edits for each (document, summary) pair.
To ensure we maintain quality control, after generating these edits, we remove the trivial edits with help of GPT4-Turbo. We ask GPT4-Turbo to classify an edit as a date change, number change, antonym change, or others. Any edit classified as date, number, or antonym change is removed from the benchmark. The prompt used in here is mentioned F.5.
The final benchmark results in 2,121 factually inconsistent summaries. To balance out the inconsistent summaries in the benchmark, we use 2,120 factually consistent edits from SummEdits benchmark resulting in a total of 4,241 samples in the final SummExecEdit benchmark. Each of the 10 domains provide around 200-300 inconsistent summaries. The distribution of each domains is given in Table 2.
Domain N %Inconsistent
SciTLDR 307 58.31
News 567 47.8
Podcast 344 58.72
BillSum 608 44.9
SamSum 450 51.11
Shakespeare 511 41.3
QMSum 334 50.3
ECTSum 416 47.12
Sales Email 368 57.1
Sales Call 336 53.87
Total 4241 50
Table 2: Statistics of the domains in SummExecEdit
@@ 5 Results
To evaluate LLMs on SummExecEdit benchmark, we use two types of prompts. D&E - Detect and Explain error. Models need to detect if there is any factual inconsistency in summary, if so, explain the inconsistency (Appendix F.3). E|D - Explain error given Detection. Given that the summary is inconsistent, models need to explain the inconsistency in the summary (Appendix F.4).
Table 4 provides the Detection Accuracy (DA) of all the models using prompt D&E. The best performing model on SummExecEdit , Claude3.5-Sonnet, provides an accuracy of only around 73%, showing challenging nature of the benchmark.
Correlation BAcc
Model V1 V2 V3 V4 V1 V2 V3 V4
claude3-haiku 0.076 0.478 0.732 0.725 0.385 0.49 0.655 0.618
claude3-opus 0.484 0.782 0.801 0.774 0.46 0.76 0.776 0.738
gpt3.5-turbo 0.103 0.118 0.325 0.563 0.375 0.401 0.495 0.625
gpt4o 0.603 0.816 0.804 0.833 0.533 0.792 0.794 0.811
5.1 Analysis for Detections
The overall detection results show that many LLMs struggle in detecting the factual error. As a reference, we also evaluate two non-LLM based approaches which use far lesser compute - AlignScore Zha et al. (2023) receiving 57.4% and MiniCheck Tang et al. (2024a) receiving 60.0% accuracy, which are better than only two open-source LLMs.
We analyze our findings and observe that 11 or more LLMs incorrectly detected nearly 1,300 out of 4,241 samples, indicating that over half of the modern day LLMs in our study struggle with more than 30% of our benchmark. Additionally, we find that 65 samples were misclassified by all LLMs. Among these, 60 samples belonged to the category of inconsistent summaries that LLMs incorrectly identified as consistent, highlighting their difficulty in detecting minor inconsistencies. Upon manually reviewing these 60 inconsistent summaries, we find that the errors primarily fall into three categories - Ungrounded information - summary contains information not grounded in the document, Reasoning Error - summary makes wrong inferences from information in the document, and Nuanced Meaning Shift - subtle shift in meaning between the document and the summary. This shows that LLMs struggle most at such cases. Examples of such misclassified cases are provided in Appendix D
@@ 5.1 Analysis for Detections
Table 4: Scores for D&E (Prompt Detect and Explain error) and E|D (Prompt Explain error given Detection) for inconsistent summaries. DA - Detection Accuracy, DS - Detection Score, ES - Explanation Score, JS - Joint Score. ES evaluated by EvalV4 prompt using GPT4o. DA based on entire benchmark. DS, ES, and JS based only on inconsistent summaries. Last 7 models are open-source.
It is worth analyzing whether newer models improve upon their predecessors in detection performance. While GPT-4o outperforms GPT-4 Turbo in detection accuracy, it struggles with explaining its reasoning, scoring lower than GPT-4 Turbo. In contrast, Claude 3.5 Sonnet and Meta’s Llama3.1 80b shows a clear improvement over Claude 3 Sonnet and Llama3 80b respectively across all aspects. A major concern arises with the Gemini Flash models. Although Gemini 2.0 Flash gets better at the explanation quality, it exhibits lower detection accuracy compared to Gemini 1.5 Flash. This raises a question if new models are truly getting better at detecting hallucination and other general purpose classification tasks. Llama3.1-8b model improves upon previous Llama3-8b but it seems that the model is more biased towards saying that the summary is inconsistent depicted from its lower detection accuracy of 0.537 but comparatively higher detection score of 0.872. Such a phenomenon is also observed for many other models.
5.2 Evaluating Explanations
The correlations for four prompts with different LLMs with respect to our manually annotated explanations are provided in Table 3. Selected explanations for manual annotations were shuffled and randomly selected. The model and either of the two prompts that generated those explanations were anonymized. The IAA between both the annotators are - Correlation of 0.885 and Cohen Kappa of 0.81. We think that the reference explanations generated at the time of edit itself are the best and of high quality, scoring 0.95 for 40 samples. The prompt EvalV4 works the best which suggests that evaluating explanations or reasoning of models, works better when we have access to what the edit/reason is or a reference explanation. Thus, good and challenging benchmarks need to also provide high quality explanations.
@@ 6 Conclusion
In this work, we first explored the executable editing with LLMs for generating benchmarks. Through our experiments, we show its superiority in generating a challenging benchmark and use it to generate a new factual inconsistency detection and explanation benchmark - SummExecEdit . We evaluate various LLMs over this benchmark for both detection and explanation of factual inconsistencies. The best performing model for detection - Claude3.7-Sonnet achieves an accuracy of 73% while Claude3-Opus performs best on combined task of detection and explanation achieving a joint score of 0.49. These scores suggest that SummExecEdit is a challenging benchmark. We explore various prompt formats that lead to better auto-evaluation by LLMs and perform error-analysis for incorrect explanations generated by LLMs. We present detection and explanation analysis that can help LLM Developers for evaluating LLMs’ abilities to detect factual errors and reason about facts before releasing the models or deploying in systems. We also hope researchers and practitioners can now use executable edits to generate datasets and benchmarks.
@@
While executable edits help in creating more challenging benchmark, it is important to note a few limitations of these edits and this work in general. First, generating benchmarks with executable edits require availability of ground truth (source, target) pairs which might not always be the case. Second, non-executable edits are equally capable of generating controlled edits but these are simple and low quality. If the end-goal is to only generate controlled edits and no-complex edits, non-executable edits could be used as well. Last, we have only experimented using these edits for creating a single benchmark related to summarization. Further experiments in other domains are required.
@@ Appendix A Model Details
Google Models. We used six Google models in our experiments - Gemini-1.5-Pro, Gemini-1.5-Flash, Gemini-Pro, Gemini-2.0 Flash, Gemini-2.0 Flash-Lite, and Palm-Bison. We used Google Cloud Platform to access Google models, which relies on the VertexAI API.
@@ Appendix B License Details and Terms of Use
Appendix C Manual Annotation and Inter-Annotator Agreement
@@
Appendix D SummExecEdit Examples
Example 1 Original Summary: A conversation between a sales agent and a potential client discussing different types of insurance products, including home, auto, life, and health insurance. The sales agent provides information on different home insurance plans and pricing, as well as available discounts for clients with good credit scores and other factors. Text to Replace: "A conversation between a sales agent and a potential client" Replaced Text: "A conversation between a sales agent and a new client" Edited Summary: A conversation between a sales agent and a new client discussing different types of insurance products, including home, auto, life, and health insurance. The sales agent provides information on different home insurance plans and pricing, as well as available discounts for clients with good credit scores and other factors. Explanation: The document indicates this is a potential client, but the summary refers to them as a new client, which is not supported.
Figure 3: Appendix D - Example 1 - SummExecEdit
Example 2 Original Summary: Abraham Lincoln Bicentennial 1-Cent Coin Redesign Act - Directs the Secretary of the Treasury, during 2009, to issue one-cent coins with the reverse side bearing four different designs representing different aspects of the life of Abraham Lincoln. Text to Replace: "during 2009" Replaced Text: "starting in 2009" Edited Summary: Abraham Lincoln Bicentennial 1-Cent Coin Redesign Act - Directs the Secretary of the Treasury, starting in 2009, to issue one-cent coins with the reverse side bearing four different designs representing different aspects of the life of Abraham Lincoln. Explanation: The document specifies that the coins are to be issued during 2009, but the summary implies they will be issued starting from 2009, potentially extending beyond that year.
Figure 4: Appendix D - Example 2 - SummExecEdit
Example 3 Original Summary: Oman is opening up its airspace to Israeli carriers, including El Al. Previously, the airlines had to fly a roundabout route to parts of Asia, adding hours to journey times. Text to Replace: "Oman is opening up its airspace to Israeli carriers, including El Al." Replaced Text: "Oman has opened up its airspace to Israeli carriers, including El Al." Edited Summary: Oman has opened up its airspace to Israeli carriers, including El Al. Previously, the airlines had to fly a roundabout route to parts of Asia, adding hours to journey times. Explanation: The document states that Oman announced it will open its airspace, implying a future event, but the summary says Oman has already opened it, suggesting it has already happened.
Figure 5: Appendix D - Example 3 - SummExecEdit
Example 4 Original Summary: Harry and Jacob discuss their preparedness for an exam that will comprise a combination of multiple choice and open question essay formats, and how knowing the format of the test can influence their study strategy. Text to Replace: "knowing the format of the test can influence their study strategy." Replaced Text: "knowing the format of the test is crucial for developing their study strategy." Edited Summary: Harry and Jacob discuss their preparedness for an exam that will comprise a combination of multiple choice and open question essay formats, and how knowing the format of the test is crucial for developing their study strategy. Explanation: While the document suggests that the test format can influence study strategy, the summary exaggerates this by stating that knowing the format is crucial for developing the strategy.
Figure 6: Appendix D - Example 4 - SummExecEdit
Example 5 Original Summary: The speaker is discussing the topic of nutrition, sharing their experience with following the Zone diet and recommending it as a way to learn about food and develop a better relationship with it. They express frustration with nutrition narratives that vilify certain foods or diets, and instead encourage focusing on the benefits of a particular diet, without putting down other approaches. Text to Replace: "and instead encourage focusing on the benefits of a particular diet, without putting down other approaches." Replaced Text: "and instead encourage focusing on the benefits of a plant-based diet, without putting down other approaches." Edited Summary: The speaker is discussing the topic of nutrition, sharing their experience with following the Zone diet and recommending it as a way to learn about food and develop a better relationship with it. They express frustration with nutrition narratives that vilify certain foods or diets, and instead encourage focusing on the benefits of a plant-based diet, without putting down other approaches. Explanation: The document mentions the speaker wants people to focus on the benefits of their chosen diet in general, but the summary specifies a plant-based diet, which is not stated in the document.
Figure 7: Appendix D - Example 5 - SummExecEdit
Appendix E Annotation Tool